OpenStudy (anonymous):

can someone remind me how to find the volume of a cone using 3.14??

OpenStudy (anonymous):

area of a cone = 1/3 pi a^2 h where h is the height and a is the area of the base

OpenStudy (anonymous):

\[1/3*\pi*r^2*h\]

OpenStudy (anonymous):

I think i tried to do the formula lenght times (1/2 *r*pi) let me try the 1/3 and see what i get thank you

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

\[\pi=3.14\]

OpenStudy (anonymous):

lol. yea i remembered that. Its been 2 years since I took college Algebra. Havent used it and have forgotten most of my formulas. howver, answers to choose from are 18 cu in 23 cu in 29 cu in 37 cu in 42 cu in When I calcualted this came up with 29.30667???

OpenStudy (anonymous):

They must be rounding the answer?? huh?

OpenStudy (anonymous):

\[1/3 *\pi*radius ^{2}* height\]